A force 1500N exists between two identical point charges separated by a distance of 30cm. Calculate the magnitude of the two point charges.


1
Expert's answer
2022-06-10T08:28:31-0400
"F=k\frac{q_1q_2}{r^2}=k\frac{q^2}{r^2}"

"q=\sqrt{Fr^2/k}\\
=\sqrt{1500*0.3^2/9*10^9}=1.2*10^{-4}\:\rm C"